OpenStudy (anonymous):

I am trying to find out the equation for the line that passes through the point (2,0) and that is parallel to the line with the equation(y-2)=1/2(x+1)

OpenStudy (anonymous):

So to find a line's equation you need to know the slope and a point. In this case you are given a point (2,0) and the line y-2=1/2(x+1). Find the slope of the equation you are given and plug that into the slope form equation which is y-y1=m(x-x1). Then you have to apply knowledge about parallel lines. If a line is parallel to another line, then its slope is its opposite reciprocal. so once you get your equation, simply take the opposite reciprocal of m. Tell me if you need further elaboration.